Q4.   Find equilibrium Price and Quantity and Price Elasticity of Demand from the following demand and supply functions

         i.       Q=10-2P                     Q=5+3P

         ii.      Qd= 200-40P                Qs= -40+80P


1
Expert's answer
2021-04-07T07:34:21-0400

i) "10-2p^{2} =5+3p"


"2p^2+3p-5=0"


"p=1, Q=8"


Price Elasticity of Demand;

"=\frac {d_Q} {d_p} ×\frac {P} {Q}"


Differentiating the demand equation we get ;

"=-4×\frac {1} {8}=-0.5"


ii) "200-40P=-40+80P"


"120P=240"

"P=2,Q=120"


Price Elasticity of Demand;we differentiate the demand equation

"\frac {d_Q} {d_P} =-40"

"=-40×\frac {2} {120}"


"=-0.6667"